# EndoUIC: Promptable Diffusion Transformer for Unified Illumination Correction in Capsule Endoscopy

### v1

@@ 4 Conclusion
This paper presents EndoUIC, a promptable DFT model for unified illumination correction for WCE. The model’s ability to navigate through different regions of the parameter space allows for tailored adjustments that address the distinct challenges posed by either overexposed or underexposed images. Furthermore, with the assistance of photographer experts, we customize the CEC dataset tailored for the EC task in WCEs. Extensive experiments conducted across four datasets demonstrate that our EndoUIC surpasses existing SOTA techniques, validating its efficacy in performing endoscopic LLIE and EC tasks. Our suggested approach can be combined with clinical endoscopy systems, greatly enhancing the visualization, diagnosis, screening, and treatment of GI diseases.
Acknowledgement
This work was supported by HK RGC GRF 14203323 & 14211420, CRF C4026-21GF, Shenzhen-HK-Macau Technology Research Programme (Type C) STIC Grant 202108233000303, and Regional Joint Fund Project 2021B1515120035 (B.02.21.00101) of Guangdong Basic and Applied Research Fund.

### v2

@@ 4 Conclusion
This paper presents EndoUIC, a promptable DiT model for unified illumination correction for WCE. The model’s ability to navigate through different regions of the parameter space allows for tailored adjustments that address the distinct challenges posed by either overexposed or underexposed images. Furthermore, with the assistance of photographer experts, we customize the CEC dataset tailored for the EC task in WCEs. Extensive experiments conducted on four datasets demonstrate that our EndoUIC surpasses existing SOTA techniques, validating its efficacy in performing endoscopic LLIE and EC tasks. Our proposed approach can be integrated with clinical endoscopy systems, greatly enhancing the visualization, diagnosis, screening, and treatment of GI diseases.
{credits}
4.0.1 Acknowledgements
This work was supported by Hong Kong RGC GRF 14211420, CRF C4063-18G, NSFC/RGC Joint Research Scheme N_CUHK420/22; Shenzhen-HK-Macau Technology Research Programme (Type C) STIC Grant 202108233000303; Regional Joint Fund Project 2021B1515120035 (B.02.21.00101) of Guangdong Basic and Applied Research Fund.
[…]
The authors have no competing interests to declare that are relevant to the content of this article.